OpenStudy (anonymous):

explain the two ways to simplify the square root of 180.and provide the simplified form.

OpenStudy (amistre64):

by hand, and by machine

OpenStudy (anonymous):

6\[\sqrt{5}\]

OpenStudy (amistre64):

1 3. 4 -------------- /1 80. 00 00 00 1 ---- 0 80 ; 2n*n < 80 : 23*3 = 69 69 ----- 11 00 ; 26n*n < 1100 : 264*4 = 1056 10 56 ------- 44 00 ; 268n*n < 4400 : etc